a ball thrown with velocity of 39.2m/s at a angle of 30` with the horizantal . calculat the maxium height and time flight
1
Expert's answer
2019-10-16T09:17:58-0400

The time flight:


"t=\frac{2v\sin{30}}{g}"

"t=\frac{2(39.2)\sin{30}}{9.8}=4\ s"

The maxium height:


"h=\frac{(v\sin{30})^2}{2g}"

"h=\frac{(39.2\sin{30})^2}{2(9.8)}=19.6\ m"